A non-dividend paying stock X is trading at $10x. Call options for 3 years maturity on the non-dividend paying stock X at a strike price of $10y and $12z cost $2t and $1x, respectively. An investor decides to enter in a bear spread position for this stock for 3 years. Calculate undiscounted profit or loss of the investor at the end of the maturity if the terminal spot price of stock X turns out to be $11t.

if you see x,y,z,t you can use it like that

x=5, y=3, z=6, t=5

Round to at least 6 decimals unless otherwise stated
Answer #1

In bar spread one purchases call with higher strike & sells call with lower strike.

Undiscounted profit=MAX(11t-12z,0)-1x-MAX(11t-10y,0)-2t=MAX(11*5-12*6,0)-1*5-MAX(11*5-10*3,0)-2*5=-40.00

Loss of $40